S how that y = c 1

e

x + c 2

e

2 x

is the general solution of y

′′ − 3 y

′ + 2y = 0 on any


interval, and find the particular solution for w hich y 0 = − 1 a n d y

( 0) = 1.


1
Expert's answer
2022-01-06T18:04:31-0500
"y''-3y'+2y=0"

Characteristic (auxiliary) equation


"r^2-3r+2=0"

"(r-1)(r-2)=0"

"r_1=1, r_2=2"

The general solution of the differential equation is


"y=c_1e^x+c_2e^{2x}, x\in \R"

Given "y(0)=-1, y'(0)=1"


"y(0)=c_1e^0+c_2e^{2(0)}"

"c_1+c_2=-1"

"y'=c_1e^x+2c_2e^{2x}"

"y'(0)=c_1e^0+2c_2e^{2(0)}"

"c_1+2c_2=1"

Then


"c_1=-1-c_2"

"c_2=2"

"c_1=-3, c_2=2"

The particular solution is


"y=-3e^x+2e^{2x}"